OpenStudy (anonymous):

what will happen to the force of gravity if the mass of one of the objects is tripled? Please help!!

OpenStudy (abhisar):

Do you know the formula for calculating gravitational force?

OpenStudy (anonymous):

F(g)=G[(m1*m2)/(r*r)]??

OpenStudy (abhisar):

Yes, now let's assume that the mass of one object each object is 5 kg and they are kept at a distance of 5 m from each other. Then the force will be \(\sf \huge \frac{G\times 5\times 5}{5\times 5}\)= G N....correct?

OpenStudy (anonymous):

yeah

OpenStudy (abhisar):

Now, let's see what happens if we triple one of the masses. Let's make one mass as 15 and keep the other one 5 as before. Now can you calculate the force for me?

OpenStudy (anonymous):

is the 15 G?

OpenStudy (anonymous):

i don't mean to seem stupid and I'm not trying to just get answers but I do have a hard time memorizing formulas

OpenStudy (abhisar):

\(\huge \frac{G\times \cancel{5}\times 15}{\cancel{5}\times5}=3G\) Getting it?

OpenStudy (abhisar):

Since the initial force was G and now after tripling one of the masses it's 3G, it seems that the force has also been tripled. Therefore we can conclude that force of gravity is directly proportional to the mass of the object.

OpenStudy (anonymous):

so if the mass of one the objects is tripled you triple the original gravitational force?

OpenStudy (abhisar):

Yeppers, and if you triple the mass of both the objects then you increase the original force with 3X3=9 times.

OpenStudy (anonymous):

okay. How will the distance between the two masses affect the gravitational forces? if you triple it the the distance increases so would it be, More separation, weaker gravitational force?

OpenStudy (abhisar):

Yes, distance is inversely proportional to the gravitational force. But there is an interesting point to be noted. If you double the distance between the two object the original force will get 1/4th because in the formula it is \(\sf Distance^2\)

OpenStudy (anonymous):

okay there's three questions in this. 1: what will happen to the force if one mass is tripled which is triple the original force 2: how will the distance between the two masses affect the gravitational forces which is more separation, weaker force and the third is The gravitational force is inversely proportional to: check all that apply... A: the product of two masses B: the universal gravitational constant C: all of these D: The square of the distance between the objects..

OpenStudy (anonymous):

I dont know the third one

OpenStudy (abhisar):

Read my last reply for getting the answer of third one :)

OpenStudy (anonymous):

I believe D is right. are there any other ones?

OpenStudy (abhisar):

Yes, D is correct. Nopes there are not any more. FYI: How to find what is directly proportional and what is indirectly proportional! Every quantity in the formula which are on the numerator side are directly proportional; and those which are on the denominator side are inversely proportional.

OpenStudy (anonymous):

can you give an example of that? like the numerator and denominator?

OpenStudy (abhisar):

Like what is numerator and what is denominator?

OpenStudy (anonymous):

how would it look in an equation? like with numbers

OpenStudy (abhisar):

Read this if you are not sure what is numerator and what is denominator http://www.bgfl.org/bgfl/custom/resources_ftp/client_ftp/ks2/maths/fractions/numerators.htm An example  Force of gravity in this case will be directly proportional to each quantity in the numerator i.e it will be directly proportional to G, mass of object one and mass of object 2. Again force will be inversely proportional to the quantity in numerator i.e square opf distance.